OpenStudy (anonymous):

What is the H+ concentration if the pH is 11.47 ?

OpenStudy (zpupster):

pH = log10[H+] [H+] = 10-pH Example: Calculate [H+] from a known pH. Find [H+] if pH = 8.5 [H+] = 10-pH [H+] = 10-8.5 [H+] = 3.2 x 10-9 M

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so would it be \[3.9 x 10^{-12}\]

OpenStudy (zpupster):

yes